OpenStudy (anonymous):

Which of the following ordered pairs is a solution to the inequality y greater than or equal to 4x - 2? (0, -4) (0, 0) (0, -3) (0, -5)

sammixboo (sammixboo):

Well let's start with choice A Plug in the values \(y \ge 4x - 2\) x = 0 y = -4 \(-4 \ge 4*0 - 2\) First multiply, what is 4*0?

OpenStudy (chosenmatt):

y greater than or equal to -2. The only pair that has an y (the 2nd number) greater or equal to -2 is the second one. also as you can see all of x is 0 so the solution will be?